# UniST: A Prompt-Empowered Universal Model for Urban Spatio-Temporal Prediction

### v4

@@ 1. Introduction
  • Extensive experiments demonstrate the generality and universality of UniST. It achieves new state-of-the-art performance on various prediction tasks, particularly, superior transferability in previously unseen scenarios.
@@ 3.3. Base Model
This transformation involves a 3D convolutional layer with a kernel size and stride both set to $(l,h,w)$. The process can be expressed as $E_{x}=\text{Conv}_{3d}(X)$, where $E_{x}$ represents the converted 1D sequential data. The sequence length of $E_{x}$ is $L^{\prime}\times H^{\prime}\times W^{\prime}$.
@@
$E_{enc}=\texttt{Encoder}(E_{x}),\ Y_{dec}=\texttt{Decoder}(E_{enc},E_{mask}),$
@@ 3.4. Spatio-Temporal Self-Supervised Pre-train
    $M=\texttt{Concat}([\textbf{1}_{(1-r)L^{\prime}\times H^{\prime}\times W^{\prime}},\textbf{0}_{rL^{\prime}\times H^{\prime}\times W^{\prime}}]),\ E_{x}=E_{x}[M=1].$
@@ 3.5.2. Spatio-Temporal Domain Knowledge
For simplicity, we provide some straightforward implementations, which are shown in the four networks in Figure 2, i.e., $Net_{tc}$, $Net_{tp}$, $Net_{sc}$, and $Net_{sh}$. For the spatial dimension, we first employ an attention mechanism to merge the temporal dimension into a representation termed $E_{s}$. Then, to capture spatial dependencies within close proximity, a 2D convolutional neural network (CNN), i.e., $Net_{sc}$, with a kernel size of 3 is employed. To capture spatial hierarchies, we utilize CNNs with larger kernel sizes, i.e., $Net_{sh}$. These larger kernels enable the perception of spatial information on larger scales, which facilitate to construct a hierarchical perspective. As for the temporal dimension, we employ an attention network, i.e., $Net_{tc}$, to aggregate the previous M steps denoted as $X_{c}$. Regarding the temporal period, we select corresponding time points from the previous N days, denoted as $X_{p}$. Subsequently, we employ another attention network, i.e., $Net_{tp}$, to aggregate the periodical sequence, which captures long-term temporal patterns. The overall process is formulated as follows:
$E_{sc}=\texttt{Conv}_{2D}[3](X_{s}),E_{sh}=\{\texttt{Conv}_{2D}[2^{i}+1](X_{s})\},i\in\{2,3,4\},$
$E_{tc}=\texttt{Attention}(X_{c}),E_{tp}=\texttt{Attention}(X_{p}).$
@@ 4.1. Experimental Setup
To evaluate the performance of UniST, we conducted extensive experiments on more than 20 spatio-temporal datasets. Due to the page limit, we select a few representative results on below, and a full benchmark can be found in the arXiv version [^1] https://arxiv.org/abs/2402.11838.
Datasets. The datasets we used cover multiple cities, spanning various domains such as crowd flow, dynamic population, traffic speed, cellular network usage, taxi trips, and bike demand. Appendix Table 4 and Table 5 (arXiv) provide a summary of the datasets we used. These spatio-temporal datasets originate from distinct domains and cities, and have variations in the number of variables, sampling frequency, spatial scale, temporal duration, and data size.
@@
Metrics. We employed commonly used regression metrics, including Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). For more detailed information of the datasets, baselines, and metrics, please refer to Appendix AB, and Appendix D (arXiv).
@@ 4.2. Short-Term Prediction
Following previous practices (Jin et al., 2023a; Nie et al., 2022), both the input step and prediction horizon are set as 6, i.e., $6\rightarrow 6$. . For baselines, we train a dedicated model for each dataset, while UniST is evaluated across all datasets.
@@ Results.
Table 2 presents the short-term prediction results, with a selection of datasets due to space constraints. The complete results can be found in Table 11 and Table 12 in Appendix E (arXiv). As we can observe from Table 2, UniST consistently outperforms all baselines across all datasets. Compared with the best baseline of each dataset, it showcases a notable average improvement. Notably, time series approaches such as PatchTST and iTransformer exhibit inferior performance compared to spatio-temporal methods. This underscores the importance of incorporating spatial dependency as prior knowledge for spatio-temporal prediction tasks. Another observation is that PatchTST(one-for-all) performs worse than PatchTST dedicated for each dataset, suggesting that the model struggles to directly adept to these distinct data distributions. Moreover, baseline approaches exhibit inconsistent performance across diverse datasets, indicating their instability across scenarios. The consistent superior performance of UniST across all scenarios underscores the significant potential and benefits of a one-for-all model. Moreover, it demonstrates UniST’s capability to orchestrate diverse data, where different datasets can benefit each other.
@@
Table 3 shows the long-term prediction results. Even with a more extended prediction horizon, UniST still consistently outperforms all baseline approaches across all datasets. Compared with the best baseline of each dataset, it yields an average improvement of 10.1%. This highlights UniST’s capability to comprehend temporal patterns effectively and its robustness in generalizing across extended durations. Table 13 in Appendix E (arXiv) illustrates the complete results.
@@ 4.4. Few-Shot Prediction
The hallmark of large foundation models lies in their exceptional generalization ability. The few-shot and zero-shot evaluations are commonly employed to characterize the ultimate tasks for universal time series forecasting (Zhou et al., 2023b; Jin et al., 2023a; Chang et al., 2023). Likewise, the few-shot and zero-shot prediction capability is crucial for a universal spatio-temporal model. In this section, we assess the few-shot learning performance of UniST. Each dataset is partitioned into three segments: training data, validation data, and test data. In few-shot learning scenarios, when confronted with an unseen dataset during the training process, we utilized a restricted amount of training data, specifically, 1%, 5%, 10% of the training data. We choose some baselines with relatively good performance for the few-shot setting evaluation, We also compare with meta-learning baselines, i.e., MAML and MetaST, and pretraining and finetuning-based time series method, i.e., PatchTST.
@@ Results.
Appendix Table 14 to 16 (arXiv) illustrates the overall few-shot results. Due to the space limit, Figure 4 only illustrates the 1% few-shot learning results on two datasets. In these cases, UniST still outperforms all baselines, it achieves a larger relative improvement over baselines compared to long-term and short-term predictions. The transferability can be attributed to successful knowledge transfer in our spatio-temporal prompt.
@@
Figure 4 also compare the performance of UniST (zero-shot) and baselines (few-shot). As observed, UniST achieves remarkable zero-shot performance, even surpassing many baselines trained with training data that are highlighted by red dashed lines. We attribute these surprising results to the powerful spatio-temporal transfer capability. It suggests that for a completely new scenario, even when the displayed overall patterns are dissimilar to the data encountered during the training process, UniST can extract fine-grained similar patterns from our defined spatial and temporal properties. The few-shot and zero-shot results demonstrate the powerful generalization capability of UniST.
5. Study and Aalysis on UniST
@@ 5.2. Prompt Learner
Next, we delve into the memorized patterns of each embedding within the temporal memory pool. Specifically, we first select the inputs based on the attention weights. For each embedding, we aggregate the corresponding input spatio-temporal data with the highest attention weight. Then, we calculate the mean value of the extracted spatio-temporal data. Figure 7(a) and Figure 7(b) illustrates the results for two datasets (Crowd and TrafficSH). As we can see, the memorized patterns revealed in the prompt tool exhibit remarkable consistency across different urban scenarios. This not only affirms that each embedding is meticulously optimized to memorize unique spatio-temporal patterns, but also underscores the robustness of the spatial and temporal memory pools across different scenarios.
Moreover, we examine the extracted spatio-temporal prompts for two distinct domains. Specifically, we calculate the mean attention weight for each embedding in the context of each dataset. Figure 7(c) and Figure 7(d) illustrates the comparison results. As we can observe, the depicted attention weight distributions for the two datasets manifest striking dissimilarities. The observed distinctiveness in attention weight distributions implies a dynamic and responsive nature in the model’s ability to tailor its focus based on the characteristics of the input data. The ability to dynamically adjust the attention weights reinforces UniST’s versatility and universality for diverse datasets.
@@ A.1. Basic Information
Here we provide more details of the used datasets in our study. We collect various spatio-temporal data from multiple cities and domains. Table 4 summarizes the basic information of the used datasets, and Table 15 (arXiv) reports the basic statistics. Specifically, values for Crowd and Cellular datasets in Table 2, Table 3, Table 13 (arXiv), Table 14 (arXiv) and Figure 4 should be scaled by a factor of $10^{3}$.
@@ Appendix B Baselines
  • ARIMA: Auto-regressive Integrated Moving Average modelis a widely used statistical method for time series forecasting. It is a powerful tool for analyzing and predicting time series data, which are observations collected at regular intervals over time.
@@
  • STGSP (Zhao et al., 2022): This model propose that the global information and positional information in the temporal dimension are important for spatio-temproal prediction. To this end, it leverages a semantic flow encoder to model the temporal relative positional signals. Besides, it utilizes an attention mechanism to cpature the multi-scale temporal dependencities.
@@
  • MAU (Chang et al., 2021): Motion-aware unit is a video predcition model. it broadens the temporal receptive fields of prediction units, which can facilitates to capture inter-frame motion correlations. It consists of an attention module and a fusion module.
@@
  • MIM (Wang et al., 2019): Memory utilize the differential information between adjacent recurrent states, which facilitates to model the non-stationary properties. Stacked multiple MIM blcoks make it possible to model high-order non-stationarity.
  • SimVP (Gao et al., 2022): It is a simple yet very effective video predcition model. It is completely built based on convulutional neural networks and uses MSE loss. It serves as a solid baseline in video prediction tasks.
  • TAU (Tan et al., 2023): Temporal Attention Unit is the state-of-the-art video predcition model. It decomposes the temporal attention into two parts: intra-frame attention and inter-frame attention, which are statical and dynamical, respectively. Besides, it introduces a novel regularization, i.e., differential divergence regularization, to consider the impact of inter-frame variations.
@@
  • STNorm (Deng et al., 2021): It proposed two types of normalization modules: spatial normalization and temporal normalization. These two normalizations can separately consider high-frequency components and local components.
@@
  • iTransformer (Liu et al., 2023a): This is the state-of-the-art multivariate time series model. Different from other Transformer-based methods, it employes the attention and feed-forward operation on an inverted dimension, that is, the multivariate correlation.
@@ E.4. Dataset Similarity
To assess the similarities among the datasets used in our study, we employed a two-step process. First, we reduced the dimensionality of the spatio-temporal data using t-SNE, a technique for dimensionality reduction. This allowed us to visualize the datasets in a lower-dimensional space. Second, we applied the k-means clustering method to the reduced data to identify clusters of similar spatio-temporal patterns.
@@ E.5. Additional Ablation Studies
E.5.2. Knowledge-Guide Prompts.
E.6. Additional Prediction Results
[…]

### v5

@@ 1. Introduction
  • Extensive experiments demonstrate the generality and universality of UniST. It achieves new state-of-the-art performance on various prediction tasks, particularly, superior few-shot and zero-shot capabilities.
@@ 3.3. Base Model
This transformation involves a 3D convolutional layer with a kernel size and stride both set to $(l,h,w)$. The process can be expressed as $E_{x}=\textsc{Conv}_{3d}(X)$, where $E_{x}$ represents the converted 1D sequential data. The sequence length of $E_{x}$ is $L^{\prime}\times H^{\prime}\times W^{\prime}$.
@@
$E_{enc}=\textsc{Encoder}(E_{x}),\ Y_{dec}=\textsc{Decoder}(E_{enc},E_{mask}),$
@@ 3.4. Spatio-Temporal Self-Supervised Pre-train
    $M=\textsc{Concat}([\textbf{1}_{(1-r)L^{\prime}\times H^{\prime}\times W^{\prime}},\textbf{0}_{rL^{\prime}\times H^{\prime}\times W^{\prime}}]),\ E_{x}=E_{x}[M=1].$
@@ 3.5.2. Spatio-Temporal Domain Knowledge
For simplicity, we provide some straightforward implementations, which are shown in the four networks in Figure 2, i.e., $\textsc{Net}_{tc}$, $\textsc{Net}_{tp}$, $\textsc{Net}_{sc}$, and $\textsc{Net}_{sh}$. For the spatial dimension, we first employ an attention mechanism to merge the temporal dimension into a representation termed $E_{s}$. Then, to capture spatial dependencies within close proximity, a two-dimensional convolutional neural network (CNN), i.e., $\textsc{Net}_{sc}$, with a kernel size of 3 is employed. To capture spatial hierarchies, we utilize CNNs with larger kernel sizes, i.e., $\textsc{Net}_{sh}$. These larger kernels enable the perception of spatial information on larger scales, which facilitate to construct a hierarchical perspective. As for the temporal dimension, we employ an attention network, i.e., $\textsc{Net}_{tc}$, to aggregate the previous M steps denoted as $X_{c}$. Regarding the temporal period, we select corresponding time points from the previous N days, denoted as $X_{p}$. Subsequently, we employ another attention network, i.e., $\textsc{Net}_{tp}$, to aggregate the periodical sequence, which captures long-term temporal patterns. The overall process is formulated as follows:
$E_{sc}=\textsc{Conv}_{2D}[3](X_{s}),$
$E_{sh}=\{\textsc{Conv}_{2D}[2^{i}+1](X_{s})\},i\in\{2,3,4\},$
$E_{tc}=\textsc{Attention}(X_{c}),$
$E_{tp}=\textsc{Attention}(X_{p}).$
@@ 4.1. Experimental Setup
To evaluate the performance of UniST, we conducted extensive experiments on more than 20 spatio-temporal datasets.
Datasets. The datasets we used cover multiple cities, spanning various domains such as crowd flow, dynamic population, traffic speed, cellular network usage, taxi trips, and bike demand. Appendix Table 4 and Table 5 provide a summary of the datasets we used. These spatio-temporal datasets originate from distinct domains and cities, and have variations in the number of variables, sampling frequency, spatial scale, temporal duration, and data size.
@@
Metrics. We employed commonly used regression metrics, including Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE). For more detailed information of the datasets, baselines, and metrics, please refer to Appendix A, Appendix B, and Appendix D.
@@ 4.2. Short-Term Prediction
Following previous practices (Jin et al., 2023a; Nie et al., 2022), both the input step and prediction horizon are set as 6, i.e., $6\rightarrow 6$. For baselines, we train a dedicated model for each dataset, while UniST is evaluated across all datasets.
@@ Results.
Table 2 presents the short-term prediction results, with a selection of datasets due to space constraints. The complete results can be found in Table 11 and Table 12 in Appendix E. As we can observe from Table 2, UniST consistently outperforms all baselines across all datasets. Compared with the best baseline of each dataset, it showcases a notable average improvement. Notably, time series approaches such as PatchTST and iTransformer exhibit inferior performance compared to spatio-temporal methods. This underscores the importance of incorporating spatial dependency as prior knowledge for spatio-temporal prediction tasks. Another observation is that PatchTST(one-for-all) performs worse than PatchTST dedicated for each dataset, suggesting that the model struggles to directly adept to these distinct data distributions. Moreover, baseline approaches exhibit inconsistent performance across diverse datasets, indicating their instability across scenarios. The consistent superior performance of UniST across all scenarios underscores the significant potential and benefits of a one-for-all model. Moreover, it demonstrates UniST’s capability to orchestrate diverse data, where different datasets can benefit each other.
@@
Table 3 shows the long-term prediction results. Even with a more extended prediction horizon, UniST still consistently outperforms all baseline approaches across all datasets. Compared with the best baseline of each dataset, it yields an average improvement of 10.1%. This highlights UniST’s capability to comprehend temporal patterns effectively and its robustness in generalizing across extended durations. Table 13 in Appendix E illustrates the complete results.
@@ 4.4. Few-Shot Prediction
The hallmark of large foundation models lies in their exceptional generalization ability. The few-shot and zero-shot evaluations are commonly employed to characterize the ultimate tasks for universal time series forecasting (Yuan et al., 2024; Zhou et al., 2023b). Likewise, the few-shot and zero-shot prediction capability is crucial for a universal spatio-temporal model. In this section, we assess the few-shot learning performance of UniST. Each dataset is partitioned into three segments: training data, validation data, and test data. In few-shot learning scenarios, when confronted with an unseen dataset during the training process, we utilized a restricted amount of training data, specifically, 1%, 5%, 10% of the training data. We choose some baselines with relatively good performance for the few-shot setting evaluation, We also compare with meta-learning baselines, i.e., MAML and MetaST, and pretraining and finetuning-based time series method, i.e., PatchTST.
@@ Results.
Appendix Table 14 to Table 16 illustrate the overall few-shot results. Due to the space limit, Figure 4 only illustrates the 1% few-shot learning results on two datasets. In these cases, UniST still outperforms all baselines, it achieves a larger relative improvement over baselines compared to long-term and short-term predictions. The transferability can be attributed to successful knowledge transfer in our spatio-temporal prompt.
@@
Figure 4 also compares the performance of UniST (zero-shot) and baselines (few-shot). As observed, UniST achieves remarkable zero-shot performance, even surpassing many baselines trained with training data that are highlighted by red dashed lines. We attribute these surprising results to the powerful spatio-temporal transfer capability. It suggests that for a completely new scenario, even when the displayed overall patterns are dissimilar to the data encountered during the training process, UniST can extract fine-grained similar patterns from our defined spatial and temporal properties. The few-shot and zero-shot results demonstrate the powerful generalization capability of UniST.
5. Study and Analysis on UniST
@@ 5.2. Prompt Learner
Next, we delve into the memorized patterns of each embedding within the temporal memory pool. Specifically, we first select the inputs based on the attention weights. For each embedding, we aggregate the corresponding input spatio-temporal data with the highest attention weight. Then, we calculate the mean value of the extracted spatio-temporal data. Figure 7(a) and Figure 7(b) illustrate the results for two datasets (Crowd and TrafficSH). As we can see, the memorized patterns revealed in the prompt tool exhibit remarkable consistency across different urban scenarios. This not only affirms that each embedding is meticulously optimized to memorize unique spatio-temporal patterns, but also underscores the robustness of the spatial and temporal memory pools across different scenarios.
Moreover, we examine the extracted spatio-temporal prompts for two distinct domains. Specifically, we calculate the mean attention weight for each embedding in the context of each dataset. Figure 7(c) and Figure 7(d) illustrate the comparison results. As we can observe, the depicted attention weight distributions for the two datasets manifest striking dissimilarities. The observed distinctiveness in attention weight distributions implies a dynamic and responsive nature in the model’s ability to tailor its focus based on the characteristics of the input data. The ability to dynamically adjust the attention weights reinforces UniST’s versatility and universality for diverse datasets.
@@ A.1. Basic Information
Here we provide more details of the used datasets in our study. We collect various spatio-temporal data from multiple cities and domains. Table 4 summarizes the basic information of the used datasets, and Table 5 reports the basic statistics. Specifically, values for Crowd and Cellular datasets in Table 2, Table 3, Table 13, Table 14 and Figure 4 should be scaled by a factor of $10^{3}$.
@@ Appendix B Baselines
  • ARIMA: Auto-regressive Integrated Moving Average model a widely used statistical method for time series forecasting. It is a powerful tool for analyzing and predicting time series data, which are observations collected at regular intervals over time.
@@
  • STGSP (Zhao et al., 2022): This model propose that the global information and positional information in the temporal dimension are important for spatio-temporal prediction. To this end, it leverages a semantic flow encoder to model the temporal relative positional signals. Besides, it utilizes an attention mechanism to capture the multi-scale temporal dependencies.
@@
  • MAU (Chang et al., 2021): Motion-aware unit is a video prediction model. it broadens the temporal receptive fields of prediction units, which can facilitates to capture inter-frame motion correlations. It consists of an attention module and a fusion module.
@@
  • MIM (Wang et al., 2019): Memory utilize the differential information between adjacent recurrent states, which facilitates to model the non-stationary properties. Stacked multiple MIM blocks make it possible to model high-order non-stationarity.
  • SimVP (Gao et al., 2022): It is a simple yet very effective video prediction model. It is completely built based on convolutional neural networks and uses MSE loss. It serves as a solid baseline in video prediction tasks.
  • TAU (Tan et al., 2023): Temporal Attention Unit is the state-of-the-art video prediction model. It decomposes the temporal attention into two parts: intra-frame attention and inter-frame attention, which are static and dynamical, respectively. Besides, it introduces a novel regularization, i.e., differential divergence regularization, to consider the impact of inter-frame variations.
@@
  • STNorm (Deng et al., 2021): It proposed two types of normalization modules: spatial normalization and temporal normalization. These two normalization methods can separately consider high-frequency components and local components.
@@
  • iTransformer (Liu et al., 2023a): This is the state-of-the-art multivariate time series model. Different from other Transformer-based methods, it employs the attention and feed-forward operation on an inverted dimension, that is, the multivariate correlation.
@@ E.4. Dataset Similarity
To assess the similarities among the datasets used in our study, we employed a two-step process. First, we reduced the dimension of the spatio-temporal data using t-SNE, a technique for dimension reduction. This allowed us to visualize the datasets in a lower-dimensional space. Second, we applied the k-means clustering method to the reduced data to identify clusters of similar spatio-temporal patterns.
@@ E.5. Additional Ablation Studies
E.5.2. Knowledge-Guide Prompts.
E.6. Additional Prediction Results
[…]